# Characterizations of voting rules based on majority margins

### v2

@@ 2.1 Preferential Equality
Let us more formally define the axiom of Preferential Equality introduced in Section 1.
@@
A voting rule $F$ satisfies Preferential Equality if for any profile $\mathbf{P}\in\mathrm{dom}(F)$ and candidates ${x,y\in X(\mathbf{P})}$, if there are two voters $i,j\in V(\mathbf{P})$ who both rank $x$ immediately above $y$, then where $\mathbf{P}^{i}$ is obtained from $\mathbf{P}$ by $i$ switching to ranking $y$ immediately above $x$, and similarly $\mathbf{P}^{j}$ is obtained from $\mathbf{P}$ by $j$ switching to ranking $y$ immediately above $x$, then $F(\mathbf{P}^{i})=F(\mathbf{P}^{j})$ (or else $\mathbf{P}^{i},\mathbf{P}^{j}$ do not belong to $\mathrm{dom}(F)$).
@@
A voting rule $F$ satisfies Preferential Compensation if for any ${\mathbf{P},\mathbf{P}^{\prime}\in\mathrm{dom}(F)}$, if $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by one voter who had $x$ immediately above $y$ in $\mathbf{P}$ switching to have $y$ immediately above $x$ in $\mathbf{P}^{\prime}$, and another voter who had $y$ immediately above $x$ in $\mathbf{P}$ switching to have $x$ immediately above $y$ in $\mathbf{P}^{\prime}$, then $F(\mathbf{P})=F(\mathbf{P}^{\prime})$.
@@
A voting rule $F$ on a rich domain of profiles satisfies Preferential Equality if and only if for any profile $\mathbf{P}$ and $n\in\mathbb{N}_{>0}$, if there are $2n$ voters who rank $x$ immediately above $y$, then for any partition of those voters into two groups $I$ and $J$ of equal size, where $\mathbf{P}^{I}$ is obtained from $\mathbf{P}$ by all voters in $I$ switching to ranking $y$ immediately above $x$, and $\mathbf{P}^{J}$ is obtained from $\mathbf{P}$ by all voters in $J$ switching to ranking $y$ immediately above $x$, we have $F(\mathbf{P}^{I})=F(\mathbf{P}^{J})$.
@@
Recall from Section 1 that Instant Runoff Voting (IRV) violates Preferential Equality, at least hypothetically. In fact, this is far from hypothetical. Table 2 shows lower bounds [^6] The reason these frequencies are only lower bounds is that we did not exhaustively search for coalitions $I$ and $J$ as described in the text but rather used a heuristic to find such coalitions $I$ and $J$. on the frequencies in four election databases of profiles $\mathbf{P}$ that witness a violation by IRV of the coalitional version of Preferential Equality in the following sense: there exist candidates $x$ and $y$ and equally sized coalitions $I$ and $J$ of voters ranking $x$ immediately above $y$ such that if, counterfactually, the voters in $I$ were to switch to ranking $y$ immediately above $x$, this would have a different effect than if instead the voters in $J$ were to make exactly the same switch. Table 2 also gives lower bounds on the frequencies of profiles $\mathbf{P}$ that witness a violation by IRV of the obvious coalitional version of Preferential Compensation, which is equivalent to Preferential Compensation: there exists candidates $x$ and $y$ and equally sized coalitions $I$ and $J$ of voters ranking $x$ immediately above $y$ and $y$ immediately above $x$, respectively, such that each coalition flipping their ranking of $\{x,y\}$ changes the IRV outcome. All the frequencies are conditional on there not being a candidate who receives a majority of first-place votes in the first round, known as an absolute majority winner (AMW), since if there is an AMW, then IRV, Plurality voting, and Condorcet voting rules all agree. The frequencies show that when IRV elections are interesting (i.e., there is no AMW), it is not unusual that there are coalitions with unequal power in terms of how they could have changed the election.
@@
# Relevant Profiles with Frequency of PEV Frequency of PCV
Dataset Profiles No AMW Conditional on No AMW Conditional on No AMW
Stable Voting Dataset 130 60 0.75 0.41
CIVS Dataset 1,216 590 0.66 0.57
PrefLib Politics (soi) 150 104 0.09 0.16
Otis 2022 dataset 448 194 0.17 0.19
Table 2: The relevant profiles are obtained from elections in which (i) every voter submits a linear order on a subset of the candidates, (ii) there is a unique IRV winner, and (iii) there are no ties when iteratively removing candidates with the fewest first-place votes; for each of these elections, we restrict the profile to the “top three” candidates, defined as the IRV winner plus the last two losers to be eliminated. The last two columns contain lower bounds on the frequencies of relevant profiles witnessing a Preferential Equality Violation (PEV) or Preferential Compensation Violation (PCV) for IRV, conditional on there not being an absolute majority winner (AMW), in elections from the Stable Voting Dataset 12-20-2024 (Holliday and Pacuit 2024), CIVS Dataset 2024-12-15 (Myers 2024a, b), PrefLib Politics (soi files as of 12-20-2024) (Mattei and Walsh 2013), and Otis 2022. Note that we restricted to elections that are not “test” elections in the CIVS Dataset.
@@ 2.2 Neutral Reversal
Next we formally define Saari’s (2003) axiom of Neutral Reversal, which has recently been used in the axiomatization of one margin-based voting rule, namely Split Cycle (see Holliday and Pacuit 2021, § 4.2, Ding et al. Forthcoming).
@@
Given a set $X\subseteq\mathcal{X}$ of candidates, a reversal pair is a pair $(L,L^{-1})$ where $L$ is a linear order of $X$ and $L^{-1}$ is the reverse of $L$ (i.e., $L^{-1}=\{(b,a)\mid(a,b)\in L\}$). We may also call a pair of voters who submit a reversal pair of rankings a “reversal pair of voters.”
A voting rule $F$ satisfies Neutral Reversal if for any $\mathbf{P},\mathbf{P}^{\prime}\in\mathrm{dom}(F)$, if $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding two voters with a reversal pair of rankings, then $F(\mathbf{P})=F(\mathbf{P}^{\prime})$.
It is easy to see that Neutral Reversal and Preferential Equality are independent axioms. Recall that the Pareto voting rule outputs for a given profile all the candidates who are not Pareto-dominated by any other candidate. Also recall that the Positive/Negative voting rule (García-Lapresta et al. 2010, Heckelman and Ragan 2021) assigns a candidate $1$ point for every voter who ranks them uniquely first and $-1$ point for every voter who ranks them uniquely last; the candidates whose sum of points is greatest are the winners. The following is straightforward to verify.
@@
    The Pareto rule, a head-to-head voting rule, satisfies Preferential Equality but not Neutral Reversal.
@@ 2.3 Proof of characterization
Given a profile $\mathbf{P}$ and $i\in V(\mathbf{P})$, let $\mathbf{P}_{i}$ be the one-voter profile that assigns to $i$ the ranking $\mathbf{P}(i)$. Given two profiles $\mathbf{P},\mathbf{Q}$ with $V(\mathbf{P})\cap V(\mathbf{Q})=\varnothing$, the disjoint union of $\mathbf{P}$ and $\mathbf{Q}$, denoted $\mathbf{P}+\mathbf{Q}$, assigns to each $i\in V(\mathbf{P})$ the ranking $\mathbf{P}(i)$ and to each $i\in V(\mathbf{Q})$ the ranking $\mathbf{Q}(i)$.
@@
Given a nonempty finite set $X\subseteq\mathcal{X}$ of candidates, a margin matrix over $X$ is a function $m:X^{2}\to\mathbb{Z}$ such that for all $x,y\in X$, $m(x,y)=-m(y,x)$. Given margin matrices, $m$ and $m^{\prime}$, we define the margin matrix $m-m^{\prime}$ by $m-m^{\prime}(x,y)=m(x,y)-m^{\prime}(x,y)$ for $x,y\in X$.
@@
[…]
@@
Let $X$ be a finite subset of $\mathcal{X}$. We define $X^{(2)}=\{(a,b)\in X^{2}\mid a\not=b\}$. For simplicity, we denote the pair $(a,b)$ simply by $ab$. For any $ab\in X^{(2)}$, with $L$ being the alphabetical order of $X\setminus\{a,b\}$, let $G^{\top}_{ab}$ be the linear order $abL$, and let $G^{\bot}_{ab}$ be the linear order $L^{-1}ab$. These two linear orders are together called a McGarvey pair.
Using this notation, we can define what is essentially Debord’s (1987) construction of profiles realizing margin graphs, generalizing McGarvey’s (1953) construction of profiles realizing majority graphs. However, for the purposes of proving Theorem 2.9, we need to be more careful about the identity of voters in the construction.
@@
Let $X$ be a nonempty finite subset of $\mathcal{X}$ and $v:X^{(2)}\times\{\top,\bot\}\times\mathbb{N}_{>0}\to\mathcal{V}$ an injective map. Under this $v$, for any $i$ in the range of $v$, when $v(ab,\star,k)=i$, we call $i$ a designated $ab$-voter and also a designated $ab$$\star$-voter.
If $m$ is a margin matrix over $X$ such that $m(x,y)$ is even for all $x,y\in X$, then we construct a profile $\mathbf{D}_{v,m}$ as follows:
@@
    $V(\mathbf{D}_{v,m})=\bigcup_{ab\in X^{(2)}}\{v(ab,\star,k)\mid 1\leq k\leq m(a,b)/2,\;\star\in\{\top,\bot\}\}$;
@@
    for any $i\in V(\mathbf{D}_{v,m})$, where $(ab,\star,k)$ is $i$’s pre-image under $v$, let $\mathbf{D}_{v,m}(i)=G^{\star}_{ab}$.
The following is easy to see from the definition of $\mathbf{D}_{v,m}$.
@@
    every voter $i\in V(\mathbf{P})$ is in the range of $v$, and where $(ab,\star,k)$ is $i$’s pre-image under $v$, we have $\mathbf{P}(i)=G_{ab}^{\star}$;
@@
    if $v(ab,\star,k)\in V(\mathbf{P})$, then for any $1\leq k^{\prime}\leq k$, we have $v(ab,\star,k^{\prime})\in V(\mathbf{P})$;
@@
[…]
@@
Suppose $\mathbf{P}$ is a profile satisfying the four conditions. By condition 1, we only need to show that $V(\mathbf{P})=V(\mathbf{D}_{v,\mathcal{M}_{\mathbf{P}}})$. For every $ab\in X^{(2)}$, let $\mathbf{P}_{|ab}$ be the profile [^7] This “profile” could be empty, but margins can be defined on the empty profile trivially. that keeps only the designated $ab$-voters in $\mathbf{P}$. Note that $\mathbf{P}$ is the disjoint union of $\mathbf{P}_{|ab}$ for all $ab\in X^{(2)}$. By conditions 1 and 2, for any $ab,cd\in X^{(2)}$ such that $\{a,b\}\neq\{c,d\}$, we have $\mathcal{M}_{\mathbf{P}_{|cd}}(a,b)=0$, since $G_{cd}^{\top}$ and $G_{cd}^{\bot}$ rank $a,b$ differently. This means $\mathcal{M}_{\mathbf{P}}(a,b)=\mathcal{M}_{\mathbf{P}_{|ab}}(a,b)-\mathcal{M}_{\mathbf{P}_{|ba}}(a,b)=|V(\mathbf{P}_{|ab})|-|V(\mathbf{P}_{|ba})|$ for any $ab\in X^{(2)}$. By condition 4, at most one of $\mathbf{P}_{|ab}$ and $\mathbf{P}_{|ba}$ can be non-empty. This means that if $\mathcal{M}_{\mathbf{P}}(a,b)\leq 0$, then $V(\mathbf{P}_{|ab})$ must be empty, so trivially we have $V(\mathbf{P}_{|ab})=\{v(ab,\star,k)\mid 1\leq k\leq\mathcal{M}_{\mathbf{P}}(a,b)/2,\star\in\{\top,\bot\}\}$. On the other hand, if $\mathcal{M}_{\mathbf{P}}(a,b)>0$, then $\mathbf{P}_{|ba}$ must be empty, meaning that $|V(\mathbf{P}_{|ab})|=\mathcal{M}_{\mathbf{P}}(a,b)$. By conditions 2 and 3, this means $V(\mathbf{P}_{|ab})$ must be $\{v(ab,\star,k)\mid 1\leq k\leq\mathcal{M}_{\mathbf{P}}(a,b)/2,\star\in\{\top,\bot\}\}$. This completes the proof. ∎
@@
It is easy to see that if $F$ is margin-based, then $F$ satisfies Preferential Equality and Neutral Reversal. Conversely, suppose $F$ satisfies Preferential Equality and Neutral Reversal. To show that $F$ is margin-based, suppose $\mathcal{M}_{\mathbf{P}}=\mathcal{M}_{\mathbf{Q}}$ (call this $m$). We must show that $F(\mathbf{P})=F(\mathbf{Q})$. Since $\mathbf{P}$ and $\mathbf{Q}$ are linear profiles with the same margins, the parity of the number of voters is the same in $\mathbf{P}$ as in $\mathbf{Q}$, and $X(\mathbf{P})=X(\mathbf{Q})$ (call it $X$). Fix an injective map
@@
Recall that if $i=v(ab,\star,k)$, $i$ is called a designated $ab$-voter and also a designated $ab$$\star$-voter.
@@
Case 1: the parity of the number of voters is even. Pick a pair of voters $i,j\in V(\mathbf{P})$. Our goal is to transform their rankings $P_{i},P_{j}$ into a reversal pair, by transforming $P_{i}$ to be the reverse of $P_{j}$. If, e.g., we want to flip $ab$ in $P_{i}$ to $ba$, pick the smallest $k\in\mathbb{N}_{>0}$ such that neither $v(ab,\top,k)$ nor $v(ab,\bot,k)$ has been added to the profile and then add them with the rankings $baL$ (not $abL$) and $L^{-1}ab$, respectively, where $L$ is the alphabetic order of $X\setminus\{a,b\}$. Thus, we have added a pair of voters with a reversal pair of rankings, which by Neutral Reversal does not change the output of $F$. Next, flip the $ab$ in $P_{i}$ to $ba$, while also changing the ranking of voter $v(ab,\top,k)$ from $baL$ to $abL$, which by Preferential Compensation (recall Lemma 2.3) does not change the output of $F$. Note that now voter $v(ab,\top,k)$’s ranking is $G^{\top}_{ab}$, and voter $v(ab,\bot,k)$’s ranking is $G^{\bot}_{ab}$. By continuing in this way, introducing sufficiently many McGarvey pairs, we transform $P_{i},P_{j}$ into the desired reversal pair. Then by Neutral Reversal, we can delete voters $i,j$ from the profile without changing the output of $F$. See Example 2.16 below for an illustration.
@@
We do this for every pair of voters from $\mathbf{P}$, obtaining a profile $\mathbf{P}^{+}$ containing none of the original voters. In fact, it is easy to observe that $\mathbf{P}^{+}$ satisfies the first three conditions of Lemma 2.15. It is also clear that $\mathcal{M}_{\mathbf{P}^{+}}=m$ since we are only using operations that maintain the margin matrix.

### v3

@@ 2.1 Preferential Equality
Let us more formally define the axiom of Preferential Equality introduced in Section 1. In the following, when we say that a voter $i$ ranks $x$ immediately above $y$, we mean that $x\mathbf{P}_{i}y$ and for all $z\in X(\mathbf{P})\setminus\{x,y\}$}, either $z\mathbf{P}_{i}x$ or $y\mathbf{P}_{i}z$.
@@
A voting rule $F$ satisfies Preferential Equality if for any profile $\mathbf{P}\in\mathrm{dom}(F)$, candidates $x,y\in X(\mathbf{P})$}, and two voters $i,j\in V(\mathbf{P})$ who both rank $x$ immediately above $y$, if $\mathbf{P}^{i}$ is obtained from $\mathbf{P}$ by $i$ switching to ranking $y$ immediately above $x$, and similarly $\mathbf{P}^{j}$ is obtained from $\mathbf{P}$ by $j$ switching to ranking $y$ immediately above $x$, then $F(\mathbf{P}^{i})=F(\mathbf{P}^{j})$ (or else $\mathbf{P}^{i},\mathbf{P}^{j}$ do not both belong to $\mathrm{dom}(F)$).
@@
A voting rule $F$ satisfies Preferential Compensation if for any $\mathbf{P},\mathbf{P}^{\prime}\in\mathrm{dom}(F)$} and $x,y\in X(\mathbf{P})$, if $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by one voter who had $x$ immediately above $y$ in $\mathbf{P}$ switching to have $y$ immediately above $x$ in $\mathbf{P}^{\prime}$, and another voter who had $y$ immediately above $x$ in $\mathbf{P}$ switching to have $x$ immediately above $y$ in $\mathbf{P}^{\prime}$, then $F(\mathbf{P})=F(\mathbf{P}^{\prime})$.
@@
A voting rule $F$ on a rich domain of profiles satisfies Preferential Equality if and only if for any profile $\mathbf{P}$, $x,y\in X(\mathbf{P})$, and $n\in\mathbb{N}_{>0}$, if there are $2n$ voters who rank $x$ immediately above $y$, then for any partition of those voters into two groups $I$ and $J$ of equal size, where $\mathbf{P}^{I}$ is obtained from $\mathbf{P}$ by all voters in $I$ switching to ranking $y$ immediately above $x$, and $\mathbf{P}^{J}$ is obtained from $\mathbf{P}$ by all voters in $J$ switching to ranking $y$ immediately above $x$, we have $F(\mathbf{P}^{I})=F(\mathbf{P}^{J})$.
@@
Recall from Section 1 that Instant Runoff Voting (IRV) violates Preferential Equality, at least hypothetically. In fact, this is far from hypothetical. Table 3 shows lower bounds [^9] The reason these frequencies are only lower bounds is that we did not exhaustively search for coalitions $I$ and $J$ as described in the text but rather used a heuristic to find such coalitions $I$ and $J$. on the frequencies in four election databases of profiles $\mathbf{P}$ that witness a violation by IRV of the coalitional version of Preferential Equality in the following sense: there exist candidates $x$ and $y$ and equally sized coalitions $I$ and $J$ of voters ranking $x$ immediately above $y$ such that if, counterfactually, the voters in $I$ were to switch to ranking $y$ immediately above $x$, this would have a different effect than if instead the voters in $J$ were to make exactly the same switch. Table 3 also gives lower bounds on the frequencies of profiles $\mathbf{P}$ that witness a violation by IRV of the obvious coalitional version of Preferential Compensation, which is equivalent to Preferential Compensation: there exists candidates $x$ and $y$ and equally sized coalitions $I$ and $J$ of voters ranking $x$ immediately above $y$ and $y$ immediately above $x$, respectively, such that each coalition flipping their ranking of $\{x,y\}$ changes the IRV outcome. We include frequencies conditional on there not being a candidate who receives a majority of first-place votes in the first round, known as an absolute majority winner (AMW), since if there is an AMW, then IRV, Plurality voting, and Condorcet voting rules all agree. The frequencies show that when IRV elections are interesting (i.e., there is no AMW), it is not unusual that there are coalitions with unequal power in terms of how they could have changed the election.
@@
Dataset Profiles Frequency of PEV Frequency of PCV
Stable Voting Dataset 216 0.35 0.22
CIVS Dataset 1,216 0.36 0.31
PrefLib Politics (soi) 308 0.03 0.05
Otis 2022 dataset 448 0.07 0.08
# Relevant Lower Bound on Lower Bound on
Profiles with Frequency of PEV Frequency of PCV
Dataset No AMW Conditional on No AMW Conditional on No AMW
Stable Voting Dataset 86 0.70 0.48
CIVS Dataset 590 0.66 0.57
PrefLib Politics (soi) 104 0.09 0.16
Otis 2022 dataset 194 0.17 0.19
Table 3: The relevant profiles are obtained from elections in which (i) each voter submits a linear order on some subset of the set of candidates (so we do not allow voters to rank candidates as tied, though we allow them to leave some candidates unranked, as in IRV elections), (ii) there is a unique IRV winner, and (iii) there are no ties when iteratively removing candidates with the fewest first-place votes; for each of these elections, we restrict the profile to the “top three” candidates (which allows the use of a simple heuristic to check for violations of Preferential Equality and Preferential Compensation), defined as the IRV winner plus the last two losers to be eliminated. The last two columns contain lower bounds on the frequencies of relevant profiles witnessing a Preferential Equality Violation (PEV) or Preferential Compensation Violation (PCV) for IRV, where in the lower half of the table these frequencies are conditional on there not being an absolute majority winner (AMW). The elections come from the Stable Voting Dataset 1-28-2026 (Holliday and Pacuit 2026), CIVS Dataset 2024-12-15 (Myers 2024a, b), PrefLib Politics (soi files as of 12-20-2024) (Mattei and Walsh 2013), and Otis 2022. Note that we restricted to elections that are not “test” elections in the CIVS Dataset.
@@ 2.2 Neutral Reversal
Next we formally define Saari’s (2003) axiom of Neutral Reversal, which has recently been used in the axiomatization of one margin-based voting rule, namely Split Cycle (see Holliday and Pacuit 2021, § 4.2, Ding et al. 2025).
@@
Given a set $X\subseteq\mathcal{X}$ of candidates, a reversal pair on $X$ is a pair $(L,L^{-1})$ where $L$ is a linear order of $X$ and $L^{-1}$ is the reverse of $L$ (i.e., $L^{-1}=\{(y,x)\mid(x,y)\in L\}$). We may also call a pair of voters who submit a reversal pair of rankings a “reversal pair of voters.”
A voting rule $F$ satisfies Neutral Reversal if for any $\mathbf{P},\mathbf{P}^{\prime}\in\mathrm{dom}(F)$, if $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding two voters with a reversal pair of rankings on $X(\mathbf{P})$, then $F(\mathbf{P})=F(\mathbf{P}^{\prime})$.
To see that Neutral Reversal and Preferential Equality are independent axioms, recall that the Positive/Negative voting rule (García-Lapresta et al. 2010, Heckelman and Ragan 2021) assigns a candidate $1$ point for every voter who ranks them uniquely first and 1-1 point for every voter who ranks them uniquely last; the candidates whose sum of points is greatest are the winners. The following is straightforward to verify.
@@
    The Pareto rule (in either its strict or weak version), a head-to-head voting rule, satisfies Preferential Equality but not Neutral Reversal.
@@ 2.3 Proof of characterization
Given a profile $\mathbf{P}$ and $i\in V(\mathbf{P})$, let $\mathbf{P}_{i}$ be the one-voter profile that assigns to $i$ the ranking $\mathbf{P}(i)$. Given two profiles $\mathbf{P},\mathbf{Q}$ with $X(\mathbf{P})=X(\mathbf{Q})$ and $V(\mathbf{P})\cap V(\mathbf{Q})=\varnothing$, the disjoint union of $\mathbf{P}$ and $\mathbf{Q}$, denoted $\mathbf{P}+\mathbf{Q}$, assigns to each $i\in V(\mathbf{P})$ the ranking $\mathbf{P}(i)$ and to each $i\in V(\mathbf{Q})$ the ranking $\mathbf{Q}(i)$.
@@
Given a nonempty finite set $X\subseteq\mathcal{X}$ of candidates, a margin matrix over $X$ is a function $m:X^{2}\to\mathbb{Z}$ such that for all $x,y\in X$, m(x,y)=m(y,x)m(x,y)=-m(y,x). Given margin matrices $m$ and $m^{\prime}$ over $X$, we define the margin matrix $m-m^{\prime}$ by $m-m^{\prime}(x,y)=m(x,y)-m^{\prime}(x,y)$ for $x,y\in X$.
@@
Given a margin matrix $m$, we want to construct a profile realizing those margins. To do so, we will use special types of rankings, specified in the following definition, also used by McGarvey (1953). First, let us fix a linear order of the set $\mathcal{X}$ of all possible candidates, which we call the alphabetic order. To avoid overloading small natural numbers with different meanings, we also fix a doubleton set $\{\top,\bot\}$.
@@
Let $X$ be a finite subset of $\mathcal{X}$. We define $X^{(2)}=\{(x,y)\in X^{2}\mid x\not=y\}$. For simplicity, we denote the pair $(x,y)$ simply by $xy$. For any $xy\in X^{(2)}$, with $L$ being the alphabetical order of $X\setminus\{x,y\}$, let $G^{\top}_{xy}$ be the linear order $xyL$, and let $G^{\bot}_{xy}$ be the linear order $L^{-1}xy$. These two linear orders are together called a McGarvey pair.
Using this notation, we can define what is essentially Debord’s (1987) construction of profiles realizing margin matrices, generalizing McGarvey’s (1953) construction of profiles realizing majority graphs. The idea for margin matrices consisting of even entries is simple: focusing on the positive entries of the margin matrix, we can uniquely decompose it as a sum of simple margin matrices with a single positive entry $2$ (and a corresponding negative entry 2-2 across the diagonal), and each of these simple margin matrices can be realized by a pair of ballots $G^{\top}_{xy}$ and $G^{\bot}_{xy}$ where $(x,y)$ is the coordinate of the entry $2$. However, for the purposes of proving Theorem 2.9—in particular in light of the absence of the axiom of anonymity in its statement—we must be able to control precisely the identity of voters in this construction; it is not enough to end up with the same anonymized profile. Thus, below we introduce a function $v$ that essentially sets up an infinite queue of voters for each type of ballot we might need. Then a Debord-style construction returns a unique profile $\mathbf{D}_{v,m}$ for each margin matrix $m$ of even parity, if we always take voters from the beginning of each queue. This $\mathbf{D}_{v,m}$ is essentially a normal form of $m$, and the main observation leading to Theorem 2.9 is that each profile $\mathbf{P}$ can be transformed into its normal form $\mathbf{D}_{v,\mathcal{M}_{\mathbf{P}}}$ using only the operations of adding or removing reversal pairs (using the voters in the queue given by $v$) and doing preferential compensation. The function $v$ will be chosen so that all voters involved are “fresh”.
@@
Let $X$ be a nonempty finite subset of $\mathcal{X}$ and $v:X^{(2)}\times\{\top,\bot\}\times\mathbb{N}_{>0}\to\mathcal{V}$ an injective map. Under this $v$, for any $i$ in the range of $v$, when $v(ab,\star,k)=i$, we call $i$ a designated $ab$-voter and also a designated abab-$\star$-voter.
If $m$ is a margin matrix over $X$ such that $m(x,y)$ is even for all $x,y\in X$ and $m$ is not the zero matrix, then we construct a profile $\mathbf{D}_{v,m}$ as follows:
@@
    $V(\mathbf{D}_{v,m})=\bigcup_{xy\in X^{(2)}}\{v(xy,\star,k)\mid 1\leq k\leq m(x,y)/2,\;\star\in\{\top,\bot\}\}$;
@@
    for any $i\in V(\mathbf{D}_{v,m})$, where $(xy,\star,k)$ is $i$’s pre-image under $v$, let $\mathbf{D}_{v,m}(i)=G^{\star}_{xy}$.
One might worry about how to define $\mathbf{D}_{v,m}$ when $m$ is a zero matrix, since by our definition all profiles must have at least one voter. In our later proof, we will introduce an independently fixed profile in this case. Now the following is easy to see from the definition of $\mathbf{D}_{v,m}$.
@@
    every voter $i\in V(\mathbf{P})$ is in the range of $v$, and where $(xy,\star,k)$ is $i$’s pre-image under $v$, we have $\mathbf{P}(i)=G_{xy}^{\star}$;
@@
    if $v(xy,\star,k)\in V(\mathbf{P})$, then for any $1\leq k^{\prime}\leq k$, we have $v(xy,\star,k^{\prime})\in V(\mathbf{P})$;
@@
    if $v(xy,\star,k)\in V(\mathbf{P})$ for some $\star\in\{\top,\bot\}$ and $k\in\mathbb{N}_{>0}$, then $v(yx,\star^{\prime},k^{\prime})\not\in V(\mathbf{P})$ for every $\star^{\prime}\in\{\top,\bot\}$ and $k^{\prime}\in\mathbb{N}_{>0}$.
@@
Suppose $\mathbf{P}$ is a profile satisfying the four conditions. By condition 1, we only need to show that $V(\mathbf{P})=V(\mathbf{D}_{v,\mathcal{M}_{\mathbf{P}}})$. For every $xy\in X^{(2)}$, let $\mathbf{P}_{|xy}$ be the profile [^10] This “profile” could be empty, but margins can be defined on the empty profile trivially. that keeps only the designated $xy$-voters in $\mathbf{P}$. Note that $\mathbf{P}$ is the disjoint union of $\mathbf{P}_{|xy}$ for all $xy\in X^{(2)}$. By conditions 1 and 2, for any $xy,x^{\prime}y^{\prime}\in X^{(2)}$ such that $\{x,y\}\neq\{x^{\prime},y^{\prime}\}$, we have $\mathcal{M}_{\mathbf{P}_{|x^{\prime}y^{\prime}}}(x,y)=0$, since $G_{x^{\prime}y^{\prime}}^{\top}$ and $G_{x^{\prime}y^{\prime}}^{\bot}$ rank $x,y$ differently. This means $\mathcal{M}_{\mathbf{P}}(x,y)=\mathcal{M}_{\mathbf{P}_{|xy}}(x,y)-\mathcal{M}_{\mathbf{P}_{|yx}}(y,x)=|V(\mathbf{P}_{|xy})|-|V(\mathbf{P}_{|yx})|$ for any $xy\in X^{(2)}$. By condition 4, at most one of $\mathbf{P}_{|xy}$ and $\mathbf{P}_{|yx}$ can be non-empty. This means that if $\mathcal{M}_{\mathbf{P}}(x,y)\leq 0$, then $V(\mathbf{P}_{|xy})$ must be empty, so trivially we have $V(\mathbf{P}_{|xy})=\{v(xy,\star,k)\mid 1\leq k\leq\mathcal{M}_{\mathbf{P}}(x,y)/2,\star\in\{\top,\bot\}\}$. On the other hand, if $\mathcal{M}_{\mathbf{P}}(x,y)>0$, then $\mathbf{P}_{|yx}$ must be empty, meaning that $|V(\mathbf{P}_{|xy})|=\mathcal{M}_{\mathbf{P}}(x,y)$. By conditions 2 and 3, this means $V(\mathbf{P}_{|xy})$ must be $\{v(xy,\star,k)\mid 1\leq k\leq\mathcal{M}_{\mathbf{P}}(x,y)/2,\star\in\{\top,\bot\}\}$. This completes the proof. ∎
@@
It is easy to see that if $F$ is margin-based, then $F$ satisfies Preferential Equality and Neutral Reversal. Conversely, suppose $F$ satisfies Preferential Equality and Neutral Reversal. To show that $F$ is margin-based, suppose $\mathcal{M}_{\mathbf{P}}=\mathcal{M}_{\mathbf{Q}}$ (call this $m$). We must show that $F(\mathbf{P})=F(\mathbf{Q})$. Since $\mathbf{P}$ and $\mathbf{Q}$ are linear profiles with the same margins (and at least $2$ candidates), the parity of the number of voters is the same in $\mathbf{P}$ as in $\mathbf{Q}$, and $X(\mathbf{P})=X(\mathbf{Q})$ (call it $X$). Fix an injective map
@@
Recall that if $v(xy,\star,k), ii$ is called a designated $xy$-voter and also a designated xyxy-$\star$-voter. Also, fix a profile with two voters not in $V(\mathbf{P})\cup V(\mathbf{Q})$ voting a reversal pair of rankings. For notational convenience, call this profile $\mathbf{D}_{v,0}$, where 0 is the zero matrix of the same dimension as $m$.
@@
Case 1: the parity of the number of voters is even. Pick a pair of voters $i,j\in V(\mathbf{P})$. Our goal is to transform their rankings $\mathbf{P}_{i},\mathbf{P}_{j}$ into a reversal pair, by transforming $\mathbf{P}_{i}$ to be the reverse of $\mathbf{P}_{j}$. If, e.g., we want to flip some adjacent candidates $xy$ in $\mathbf{P}_{i}$ to $yx$, pick the smallest $k\in\mathbb{N}_{>0}$ such that neither $v(xy,\top,k)$ nor $v(xy,\bot,k)$ has been added to the profile and then add them with the rankings $yxL$ (not $xyL$) and $L^{-1}xy$, respectively, where $L$ is the alphabetic order of $X\setminus\{x,y\}$. Thus, we have added a pair of voters with a reversal pair of rankings, which by Neutral Reversal does not change the output of $F$. Next, flip the $xy$ in $\mathbf{P}_{i}$ to $yx$, while also changing the ranking of voter $v(xy,\top,k)$ from $yxL$ to $xyL$, which by Preferential Compensation (recall Lemma 2.3) does not change the output of $F$. Note that now voter $v(xy,\top,k)$’s ranking is $G^{\top}_{xy}$, and voter $v(xy,\bot,k)$’s ranking is $G^{\bot}_{xy}$. By continuing in this way, introducing sufficiently many McGarvey pairs, we transform $\mathbf{P}_{i},\mathbf{P}_{j}$ into the desired reversal pair. Then by Neutral Reversal, we can delete voters $i,j$ from the profile without changing the output of $F$. See Example 2.16 below for an illustration.
@@
We do this for every pair of voters from $\mathbf{P}$, obtaining a profile $\mathbf{P}^{+}$ containing none of the original voters. In fact, it is easy to observe that $\mathbf{P}^{+}$ satisfies the first three conditions of Lemma 2.15, since all old voters in $\mathbf{P}$ have been deleted, and new voters were taken from the front of the each queue given by $v$ in pairs. It is also clear that $\mathcal{M}_{\mathbf{P}^{+}}=m$ since we are only using operations that maintain the margin matrix.